OpenStudy (anonymous):

Condense 1/2 log516 - 3 log5X + 4 log5Y

OpenStudy (experimentx):

log((516)^1/2) - log((5X)^3) + log((4Y)^4) log((516)^1/2*(4Y)^4/(5X)^3)

OpenStudy (anonymous):

1/2log51693log5x+log5y let a=1/2log516 b=3log5x c=4log5y then a-b+c a/bc